Suppose you have $1000 invested in a fund that pays 6% per year compounded monthly. a) Write an equation for the function that fives the value of the investment after "m" months. I came up with V=1000(1+0.06 divided by 12)^m
b) Then, Find an equation for the inverse of the function found in part (a).

Found 2 solutions by stanbon, Boreal:
Answer by stanbon(75887)

Suppose you have $1000 invested in a fund that pays 6% per year compounded monthly.
a) Write an equation for the function that finds the value of the investment after "m" months. I came up with V=1000(1+0.06 divided by 12)^m
V(t) = 1000(1.005)^t
-------------------------
b) Then, Find an equation for the inverse of the function found in part (a).
Interchange V and t to get:
t = 1000(1.005)^V
------
Solve for "V"::
1.005^V = t/1000
V = log(t/1000)/log(1.005)
====================
Cheers,
Stan H.
-------------------

Answer by Boreal(15235)

You are correct:
P=1000[1 + (0.06/12)]^m
P=1000(1.005)^m
Isolate m to get the inverse.
{P/1000) = (1.005)^m
log (P/1000)= m log (1.005)
Taking the log removes the exponent.
m=log (P/1000)/0.00217 the denominator is the log (1.005)
OR
m=(log P - log 1000)/.00217
m=(log P -3)/0.00217